The width of a rectangle is 2 m less than half the length. The perimeter is 44 m.Find the dimensions.
Answer by checkley77(12569)

W=L/2-2
2L+2W=44
2L+2(L/2-2)=44
2L+L-4=44
3L=44+4
3L=48
L=48/3
L=16 ANS. FOR THE LENGTH.
W=16/2-2
W=8-2
L=6 ANS. FOR THE WIDTH.
PROOF:
2*16+2*6=44
32+12=44
44=44